1 18) What is the value of K, for this reaction at 823 K. Fe,03 (s) + 3CO (g) ? 2 Fe (s) + 3CO2 (g) K,=67 @ 823 K A) 0.99 B 1.01 C)67 (D)y.5 x 10 (2
Answer #1

We know that Kp and Kc are related for a gaseous reaction as

Kp = Kc*(RT)?n

where ?n = change in number of moles of gaseous species.

We have n = 3 moles of gaseous reactants on the right and n = 3 moles of gaseous products on the right. Therefore, we have ?n = 3 – 3 = 0 and (RT)?n = 0. Thus,

Kp = Kc*(RT)?n = (67)*(RT)0 = (67)*(1) = 67.

(C) is the correct answer.